OpenStudy (anonymous):

X^2+2y=9 x-y+3=0 could someone show me the steps to solving this?

OpenStudy (anonymous):

If you are finding x and y then : 1) call the first equation(a) and second equation(b) 2) solve for y in equation(b) and you'll get: \[y = x + 3\] then substitute y in equation(a) and you'll get : \[x^2 + 2(x + 3) - 9 = 0 \] \[x^2 + 2x -3 = 0\]\[(x-1)(x+3) = 0\] then you have : x = 1 and x = -3 Substitute both x's in equation (b) to find y, so for the first x = 1, you'll get : y = 1 + 3 y = 4 and for the second x = -3, you'll get : y = -3 + 3 y = 0 In summary, your 2 points are : (1,4) and (-3,0) ^_^

OpenStudy (anonymous):

do you understand how to solve such problems now? :)

OpenStudy (anonymous):

absolutly! thanks!